# The Completeness of Eigenstates in Quantum Mechanics

### v1

@@ IV.1.1 The set of eigenstates in unequal constant potential
$\psi^{0}(x)=\begin{cases}\mathcal{A}e^{ikx}+\mathcal{B}e^{-ikx},&x<0\\ \mathcal{C}e^{-\alpha x},&x>0\end{cases}.$ (47)
@@
$\psi(x)=\begin{cases}\mathcal{D}\cos(kx)+\mathcal{E}\sin(kx),&x<0\\ \mathcal{F}\cos(\beta x)+\mathcal{G}\sin(\beta x),&x>0\end{cases}.$ (52)
@@
${\psi}^{1}(x)=\begin{cases}\mathcal{F}\cos(kx)+\sqrt{\frac{2\beta^{2}\left(1-\pi\mathcal{F}^{2}\right)}{\pi\left(\beta^{2}+k^{2}\right)}}e^{i\theta}\sin(kx),&x<0\\ \mathcal{F}\cos(\beta x)+\sqrt{\frac{2k^{2}\left(1-\pi\mathcal{F}^{2}\right)}{\pi\left(\beta^{2}+k^{2}\right)}}e^{i\theta}\sin(\beta x),&x>0\end{cases};$ (65)
${\psi}^{2}(x)=\begin{cases}\sqrt{\frac{1-\pi\mathcal{F}^{2}}{\pi}}\cos(kx)-\sqrt{\frac{2\beta^{2}\mathcal{F}^{2}}{\beta^{2}+k^{2}}}e^{i\theta}\sin(kx),&x<0\\ \sqrt{\frac{1-\pi\mathcal{F}^{2}}{\pi}}\cos(\beta x)-\sqrt{\frac{2k^{2}\mathcal{F}^{2}}{\beta^{2}+k^{2}}}e^{i\theta}\sin(\beta x),&x>0\end{cases}.$ (66)
@@
${\psi}^{0}(x)=\begin{cases}\sqrt{\frac{2k^{2}}{\pi\left(\alpha^{2}+k^{2}\right)}}\cos(kx)-\sqrt{\frac{2\alpha^{2}}{\pi\left(\alpha^{2}+k^{2}\right)}}\sin(kx),&x<0\\ \sqrt{\frac{2k^{2}}{\pi\left(\alpha^{2}+k^{2}\right)}}e^{-\alpha x},&x>0\end{cases}.$ (67)
@@ IV.1.2 The spectral function $\kappa(\varepsilon)$ in unequal constant potential
Suppose $\tau<0$, $\sigma>0$, and $\tau+\sigma<0$ — i.e., the initial state is entirely confined to the region $x<0$. Next, fixing the initial state width $\sigma$ and taking the limit $\tau\to$-$\infty$ in Eq. 70, we apply the Riemann-Lebesgue lemma [Riemann] to obtain the following equation:
@@
${\psi}^{1}(x)=\begin{cases}\frac{1}{\sqrt{\pi}}\cos(kx),&x<0\\ \frac{1}{\sqrt{\pi}}\cos(\beta x),&x>0\end{cases};$ (84)
${\psi}^{2}(x)=\begin{cases}\sqrt{\frac{2\beta^{2}}{\pi\left(\beta^{2}+k^{2}\right)}}\sin(kx),&x<0\\ \sqrt{\frac{2k^{2}}{\pi\left(\beta^{2}+k^{2}\right)}}\sin(\beta x),&x>0\end{cases}.$ (85)
@@ IV.1.3 The completenes of the set of eigenstates in unequal constant potential
Complexifying the integrand by substituting $k\to z$, we note that in the upper half-plane, $\frac{\pi}{2}<\frac{1}{2}\left(\lambda(z)+\rho(z)\right)<\frac{3\pi}{2}$, which satisfies the conditions of Jordan’s lemma [Jordan]. Integrating along the contour illustrated in Fig. 5c and applying the residue theorem [Jordan], we obtain the following result:
@@
The completeness proofs for the regions $x<0$ and $x>\sigma$ are analogous to this and are omitted here to avoid redundancy. The proofs for the points $x=0$ and $\sigma$ follow from the continuity of an improper integral depending on a parameter[Zorich1]. Specifically, the integral in Eq. 87 at a fixed point $x$ is equal to the limit of Eq. 92 at the same fixed point $x$. Since the initial state is a continuous function, the completeness at the points $x=0$ and $\sigma$ is thus established.
@@
$F(x)=\begin{cases}\phi(x),&-\chi<x<0\\ 0,&0\leq x<\chi\end{cases};$
$G(x)=\begin{cases}0,&-\chi<x\leq 0\\ \phi(x),&0<x<\chi\end{cases}.$
@@
${\phi}_{i}(x)=\begin{cases}\sqrt{\frac{2}{\chi}}\sin\left(\frac{i\pi(x+\chi)}{\chi}\right),&-\chi<x<0\\ 0,&\text{otherwise}\end{cases};$
${\phi}_{j}(x)=\begin{cases}0,&\text{otherwise}\\ \sqrt{\frac{2}{\chi}}\sin\left(\frac{j\pi(x+\chi)}{\chi}\right),&0<x<\chi\end{cases}.$
@@
At $x=0$, since the initial state $\phi(x)$ and eigenstates $\psi_{\kappa}(x)$ are continuous, it follows from the continuity of an integral depending on a parameter[Zorich2] that:
@@ IV.2.1 The set of eigenstates in one-sided open box potential
$V(x)=\begin{cases}+\infty,&x<0\\ -V_{0},&0<x<1\\ 0,&x>1\end{cases}.$ (95)
@@
${\psi}_{n}(x)=A\begin{cases}\sin(\beta x),&0<x<1\\ e^{k}\sin\left(\beta\right)e^{-kx},&x>1\end{cases}.$ (97)
@@
${\psi}_{\kappa}(x)=\begin{cases}\frac{\kappa\sin(\gamma x)}{\sqrt{\pi\left(\gamma^{2}\cos^{2}{\gamma}+\kappa^{2}\sin^{2}{\gamma}\right)/2}},&0<x<1\\ \frac{\kappa\sin\gamma\cos(\kappa x-\kappa)+\gamma\cos\gamma\sin(\kappa x-\kappa)}{\sqrt{\pi\left(\gamma^{2}\cos^{2}{\gamma}+\kappa^{2}\sin^{2}{\gamma}\right)/2}},&x>1\end{cases}.$ (98)
@@ IV.2.2 The completeness of the set of eigenstates in one-sided open box potential
$\mathcal{V}\left(x\right)=\begin{cases}+\infty,&x<0\\ -\mathcal{V}_{0},&0<x<2/3\\ -\mathcal{V}_{0}+3\mathcal{V}_{0}(x-2/3),&2/3<x<1\\ 0,&x>1\end{cases}.$
@@ IV.3.2 Orthonormalization of Eigenstates
$1=\underset{m\to+\infty}{\text{lim}}\frac{\int_{0}^{ma}\left(\frac{1}{\sqrt{\pi}}\right)^{2}\,dx}{\sum_{n=1}^{m}\left(|A_{n}|^{2}+|B_{n}|^{2}\right)a}$ (111)
@@ IV.4 Finite periodic potential on one side
$V_{1}(x)=\begin{cases}+\infty,&x<0\\ \sin(2x),&x>0\end{cases}.$
@@
$V_{4}(x)=\begin{cases}-\mathcal{V}_{0}\cos(\pi x/a),&x<0\\ \mathcal{V}_{1}\cos(\pi x/b),&x>0\end{cases}.$
@@
As for the spectral function $\kappa(\varepsilon)$ used for eigenstate expansion, its determination is even more challenging. Notably, while general mathematical methods for determining the energy spectrum of periodic potentials exist in the literature [Arscott][Frank], the determination of the energy spectrum for one-sided periodic potentials is established herein through our work.

### v2

@@ IV.1.1 The set of eigenstates in unequal constant potential
$\psi^{0}(x)=\cases{}{A}e^{ikx}+\mathcal{B}e^{-ikx},&x<0\\ \mathcal{C}e^{-\alpha x},&x>0.$ (47)
@@
$\psi(x)=\cases{}{D}\cos(kx)+\mathcal{E}\sin(kx),&x<0\\ \mathcal{F}\cos(\beta x)+\mathcal{G}\sin(\beta x),&x>0.$ (52)
@@
${\psi}^{1}(x)=\cases{}{F}\cos(kx)+\sqrt{\frac{2\beta^{2}\left(1-\pi\mathcal{F}^{2}\right)}{\pi\left(\beta^{2}+k^{2}\right)}}e^{i\theta}\sin(kx),&x<0\\ \mathcal{F}\cos(\beta x)+\sqrt{\frac{2k^{2}\left(1-\pi\mathcal{F}^{2}\right)}{\pi\left(\beta^{2}+k^{2}\right)}}e^{i\theta}\sin(\beta x),&x>0;$ (65)
${\psi}^{2}(x)=\cases{\sqrt{\hfil}}{\frac{1-\pi\mathcal{F}^{2}}{\pi}}\cos(kx)-\sqrt{\frac{2\beta^{2}\mathcal{F}^{2}}{\beta^{2}+k^{2}}}e^{i\theta}\sin(kx),&x<0\\ \sqrt{\frac{1-\pi\mathcal{F}^{2}}{\pi}}\cos(\beta x)-\sqrt{\frac{2k^{2}\mathcal{F}^{2}}{\beta^{2}+k^{2}}}e^{i\theta}\sin(\beta x),&x>0.$ (66)
@@
${\psi}^{0}(x)=\cases{\sqrt{\hfil}}{\frac{2k^{2}}{\pi\left(\alpha^{2}+k^{2}\right)}}\cos(kx)-\sqrt{\frac{2\alpha^{2}}{\pi\left(\alpha^{2}+k^{2}\right)}}\sin(kx),&x<0\\ \sqrt{\frac{2k^{2}}{\pi\left(\alpha^{2}+k^{2}\right)}}e^{-\alpha x},&x>0.$ (67)
@@ IV.1.2 The spectral function $\kappa(\varepsilon)$ in unequal constant potential
Suppose $\tau<0$, $\sigma>0$, and $\tau+\sigma<0$ — i.e., the initial state is entirely confined to the region $x<0$. Next, fixing the initial state width $\sigma$ and taking the limit $\tau\to$-$\infty$ in Eq. 70, we apply the Riemann-Lebesgue lemma [2] to obtain the following equation:
@@
${\psi}^{1}(x)=\cases{}{1}{\sqrt{\pi}}\cos(kx),&x<0\\ \frac{1}{\sqrt{\pi}}\cos(\beta x),&x>0;$ (84)
${\psi}^{2}(x)=\cases{\sqrt{\hfil}}{\frac{2\beta^{2}}{\pi\left(\beta^{2}+k^{2}\right)}}\sin(kx),&x<0\\ \sqrt{\frac{2k^{2}}{\pi\left(\beta^{2}+k^{2}\right)}}\sin(\beta x),&x>0.$ (85)
@@ IV.1.3 The completenes of the set of eigenstates in unequal constant potential
Complexifying the integrand by substituting $k\to z$, we note that in the upper half-plane, $\frac{\pi}{2}<\frac{1}{2}\left(\lambda(z)+\rho(z)\right)<\frac{3\pi}{2}$, which satisfies the conditions of Jordan’s lemma [1]. Integrating along the contour illustrated in Fig. 5c and applying the residue theorem [1], we obtain the following result:
@@
The completeness proofs for the regions $x<0$ and $x>\sigma$ are analogous to this and are omitted here to avoid redundancy. The proofs for the points $x=0$ and $\sigma$ follow from the continuity of an improper integral depending on a parameter[31]. Specifically, the integral in Eq. 87 at a fixed point $x$ is equal to the limit of Eq. 92 at the same fixed point $x$. Since the initial state is a continuous function, the completeness at the points $x=0$ and $\sigma$ is thus established.
@@
$F(x)=\cases{\phi}(x),&-\chi<x<0\\ 0,&0\leq x<\chi;$
$G(x)=\cases{0},&-\chi<x\leq 0\\ \phi(x),&0<x<\chi.$
@@
${\phi}_{i}(x)=\cases{\sqrt{\hfil}}{\frac{2}{\chi}}\sin\left(\frac{i\pi(x+\chi)}{\chi}\right),&-\chi<x<0\\ 0,&\text{otherwise};$
ϕj(x)={0,otherwise
2χsin(jπ(x+$χ)$
,0<x<χ
[…]
@@
At $x=0$, since the initial state $\phi(x)$ and eigenstates $\psi_{\kappa}(x)$ are continuous, it follows from the continuity of an integral depending on a parameter[32] that:
@@ IV.2.1 The set of eigenstates in one-sided open box potential
$V(x)=\cases{+}\infty,&x<0\\ -V_{0},&0<x<1\\ 0,&x>1.$ (95)
@@
${\psi}_{n}(x)=A\cases{\sin}(\beta x),&0<x<1\\ e^{k}\sin\left(\beta\right)e^{-kx},&x>1.$ (97)
@@
${\psi}_{\kappa}(x)=\cases{}{\kappa\sin(\gamma x)}{\sqrt{\pi\left(\gamma^{2}\cos^{2}{\gamma}+\kappa^{2}\sin^{2}{\gamma}\right)/2}},&0<x<1\\ \frac{\kappa\sin\gamma\cos(\kappa x-\kappa)+\gamma\cos\gamma\sin(\kappa x-\kappa)}{\sqrt{\pi\left(\gamma^{2}\cos^{2}{\gamma}+\kappa^{2}\sin^{2}{\gamma}\right)/2}},&x>1.$ (98)
@@ IV.2.2 The completeness of the set of eigenstates in one-sided open box potential
$\mathcal{V}\left(x\right)=\cases{+}\infty,&x<0\\ -\mathcal{V}_{0},&0<x<2/3\\ -\mathcal{V}_{0}+3\mathcal{V}_{0}(x-2/3),&2/3<x<1\\ 0,&x>1.$
@@ IV.3.2 Orthonormalization of Eigenstates
[…]
@@ IV.4 Finite periodic potential on one side
$V_{1}(x)=\cases{+}\infty,&x<0\\ \sin(2x),&x>0.$
@@
$V_{4}(x)=\cases{-}\mathcal{V}_{0}\cos(\pi x/a),&x<0\\ \mathcal{V}_{1}\cos(\pi x/b),&x>0.$
@@
As for the spectral function $\kappa(\varepsilon)$ used for eigenstate expansion, its determination is even more challenging. Notably, while general mathematical methods for determining the energy spectrum of periodic potentials exist in the literature [3][21], the determination of the energy spectrum for one-sided periodic potentials is established herein through our work.